Oxford Materials IT Committee Meeting

MICHELMAS TERM 2005
A meeting of the Department of Materials IT Committee was held on Tuesday 25th October 2005.

AGENDA

  • 1 Apologies for absence and Membership of the Committee
  • 2 Matters arising from the Minutes of the Meeting of 10th May 2005
  • 3 IT/AV – Matters for Discussion (attached)
  • 4 IT/AV – Head of IT Report (attached)
  • 5 University IT Users Group – brief verbal report (PJW)
  • 6 IT/AV expenditure - matters for report (PJW)
  • 7 AOB
  • 8 Date and time of next meeting

ITEM 2. MATTERS ARISING FROM THE MEETING OF 10th MAY 2005

Department Photograph (Minute 2a)
Departmental Photograph will be taken by digital camera only. Requests for hard copies of the photograph was co-ordinated by AHM as usual and at similar cost to previous years. Electronic copy was made available to Department members via the department website.

University Anti-virus policy (Minute 5a)
IT staff contributed to the proposed university-wide Anti-virus policy under discussion.

IDSAF (Minute 7c)
AMcK raised concerns over the IDSAF proposal to centralize the management of the Media Services (photographic, AV etc), and in particular possible personnel relocation. RMP noted the concerns and briefly outlined the IDSAF project http://www.mpls.ox.ac.uk/intranet/IDSAF for the benefit of JAMS. RMP communicated these views to the project manager and to pointed out that there now needs to be wider consultation with the Department user-base. AMcK had a meeting with the project manager.

ITEM 3: IT/AV MATTERS FOR DISCUSSION - 10th MAY 2005

None.

ITEM 4: IT/AV HEAD OF IT REPORT – 10th MAY 2005

Admin Network
Several staff changes are due in the near future. (GDWS->CRMG, RMP->APD, RAM->TMcA)

Teaching Laboratory
The general scanner was replaced with a Epson Perfection 3490 Photo Scanner. The colour printer needs fixing or replacing.

Lecture theatres / meeting rooms
The CETL bid for equipment to link the HR lecture theatre to Begbroke is proceeding slowly.

Digital Darkroom
Networks problems affecting printing were resolved by swapping fibre converter cable for the building (possibly affected by building works). The scanner was replaced with an Epson Perfection 4990 Photo Scanner.

Library
No issues to report. The systems were checked this term.

MML
Lots of activity and changes in progress. A test cluster has been installed and configured and is being expanded to satisfy temporary requirements prior to the new 1000node cluster due in the OSC in May 2006. A replacement fileserver is being configured and gigabit networking installed. Compiler licenses have been made available free of charge university-wide via comlab. Re-wiring plans are progressing slowly as part of the building rewiring.

EM
The EM IT room is currently being supported by IT staff, but at a reduced level until the appointment of a replacement technician. The Cothi usage-monitoring system was installed and is working.

Website
The intended review of research group websites over the summer has not happened yet. The OCAMAC website was rewritten to be compliant with current guidelines. Most information has been updated for the start of the new academic year. Research in Progress 2005 is nearly finalized after which the research information on the website will be updated. A Materials presence on http://weblearn.ox.ac.uk has been created but is still hidden/under development.

Publications
All handbooks for 2005 have been updated and printed except for Research in Progress 2005.

Network
Finally the new 8 km optical fibre to Begbroke was completed and activated. Initial troubles were traced to intermittent signal quality issue with a 0.5m fibre patch cable which was replaced. A bridging firewall has been built and tested. The firewall and associated logfile analysis and registration procedures are still be refined.

Servers
Materials2(main netware fileserver) and Nomad (deployment, antivirus, groupshares) remain unchanged. New webservers www-int, www-ext and www-test have been installed and the old webservers databases and resolve2 are almost ready for decommissioning. The new wsus (windowsupdates) server has been installed and the old sus(windowsupdates) server will be decommissioned when all users have been transferred. The possibility of running an Exchange server is being investigated but at low priority using Materials1.

Phones
One phone was replaced. One day of connectivity problems was traced to a network switch needing restarting.

Security
No major breaches. One user reprimanded for p2p filesharing. The new Sophos Enterprise server has been installed

Begbroke Directorate
Directorate no-longer supported by Materials IT staff.

Miscellaneous
• A new A3 colour printer to be located on the top floor of Hume-Rothery has been purchased.

• The outcome of a team design project on recycling across the department is expected to impact IT procedures, especially printing.

• The system for research groups purchasing desktops from IT stock is working well and widely used. Laptops are still always configured on request.


MINUTES

A meeting of the Department of Materials IT Committee was held on Tuesday 10th May 2005.
The following people attended: PJW, AMC, CJDH, JAMS.

1. Apologies for Absence and Membership of the Committee

The committee welcomed Khalid Schofield (KS) who was appointed as an IT Technician effective from 13th June 2005. The committee noted the departure of the Department Administrator Rosie Plummer, thanked her for her contribution to the committee and wished her well for the future. Ms Alana Davies will start as the new Administrator in January 2006. Apologies were received from KS (on holiday) and AHM (videoing a lecture).

2. Matters arising from the minutes

a) Lecture theatres / meeting rooms (Minute 2b)
The University-wide CETL project, for which Adrian Taylor is the department contact, is progressing towards purchase of IT equipment. The time-scale for the provision of the video conference link between Begbroke and the Hume-Rothery Lecture theatre is still unclear but division-wide the plans are progressing. BR lecture theatre is likely to be refurbished which may affect implementation of CETL equipment. PJW to monitor progress.

Action: PJW

b) Email and shared calendar server (Minute 2c)
Requirements of an Exchange server have been investigated. Informal comments on Lotus Notes as used in Central Admin were unfavourable.

Action: AMC/PJW

c) IT disaster recovery plans (Minute 3a)
PJW to circulate draft documentation for comment and to start the process of collecting full details. Carried to next meeting.

Action: PJW

d) Gigabit networking (Minute 3b)
i) The new fibre connection was installed directly within the Materials subnet, thus precluding the need for a second backbone connection.
ii) Upgrading the existing backbone connection from 100Mbit to Gigabit was deemed unnecessary, given current network usage. Usage has been monitored and will reported at future meetings.
iii) The microwave link is to be retired (terminating the annual maintenance £7k/year). PJW to investigate the cost and technical requirements for a (low bandwidth) backup network connection as part of the disaster recovery planning.

Action: PJW

iv) Also discussed was the possible effect of upgrading the network to Gigabit upon the Quality of Service for connection sensitive applications such as VOIP phones and videoconferencing. PJW to investigate.

Action: PJW

e) Begbroke Directorate (Minute 4a)
RMP contacted the Begbroke Directorate to discuss arrangements for sharing IT support provision at Begbroke to cover holidays and illness etc.

Action: PJW

f) Website (Minute 4b)
The proposed review of research group websites did not happen during the summer but has been carried forward to the next meeting. Advice and guidance is to be offered by PJW upon compliance with University Website Policies and upon mechanisms for easy updating of content. PJW to email notices to raise interest in web authoring and best practice.

Action: PJW

3 Matters for Discussion

None.

4. Head of IT Report

A written report was circulated with the agenda (above).

5. Oxford University IT Users Group Meeting http://www.ict.ox.ac.uk/oxford/groups/ouitug/

a) ICT

A high level ICT strategy workgroup has been convened to generate an IT strategy for the whole university by April 2006. At the same time a small ICT transition team is working towards common IT provision across Central Administration, OUCS and OULS. Major strategic organisational changes are being considered. Department IT staff to monitor progress. http://www.ict.ox.ac.uk/strategy/

b) IT Metrics

The first part of the IT Metrics project was completed and results of the survey of IT infrastructure/procedures were collated and circulated (winning a nomination for incomprehensibility). The second part of the IT Metrics project surveying the IT skills provision across the university is due to start in February 2006.

6. IT/AV Expenditure

Figures were not available.

7 AOB

a) Membership
Crispin Hetherington resigned from the ITC committee. Crispin was thanked for his contribution to the Committee. PJW to forward nominations for his replacement to DC.

Action: PJW

b) Terms of Reference
Changes to the terms of reference http://www.materials.ox.ac.uk/it/itcommittee.html as follows:
- 2 members of Department nominated annually bi-annually by DC
- The committee shall meet at least once per term, usually on Monday of 1st Week Tuesday of 3rd Week.

8 Date of Next Meeting

The next meeting will take place on Tuesday 24th January 2006.

The meeting closed at 2.45pm